# Structure-Accurate Medical Image Translation based on Dynamic Frequency Balance and Knowledge Guidance

### v1

@@ 1 Introduction
To further improve image generation quality, researchers begin to introduce the prior knowledge of vision language model (VLM) into the generation process [21, 22, 23]. The pre-trained VLM based on the billion-parameter scale provides rich prior knowledge for image translation, greatly reducing the learning difficulty while ensuring that the generated images have the correct anatomical structure. However, existing algorithms often blindly accept all prior knowledge, resulting in a failure to effectively fuse prior knowledge with the model, which has become a challenge in image translation [24].
@@ 3.2 Medical Image Translation with DFBK
$z^{L}_{s}=S(\frac{Q_{z^{ll}}K_{z^{ll}}^{T}}{\sqrt{d}})V_{z^{ll}},$ (5)
In order to achieve dynamic frequency balance, we use the differential attention [40] mechanism for high-frequency features,
where $\lambda_{q}^{1}$, $\lambda_{q}^{2}$, $\lambda_{k}^{1}$, $\lambda_{k}^{2}$ are learnable parameters, and $\lambda_{init}$ is a constant. Differential attention can dynamically filter out and enhance effective high-frequency information while suppressing useless high-frequency noise.
$z_{s}=IWT(z^{L}_{s},z^{H}_{s}),$ (8)
$z^{L}_{k}=S(\frac{Q_{z^{ll}}K_{c}^{T}}{\sqrt{d}})V_{c}^{1},$ (9)
Finally, we can obtain the recombination feature $z_{k}$,
$z_{k}=IWT(z^{L}_{k},z^{H}_{k}).$ (11)
@@ 4 Implementation Details
Figure 3: Comparison of approaches on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated image and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa (e.g., blue).
Figure 4: Comparison of approaches on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 5: Comparison of approaches on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of CT to MR and MR to CT translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated image and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Table 1: Quantitative comparison of different approaches on BraTs and ISLES. The values in the table represent the average values for the entire test dataset. Bold represents the best result and underlined represents the suboptimal result.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
Ea-GAN 22.43 0.825 0.083 21.16 0.773 0.091 24.18 0.850 0.065 21.35 0.810 0.092
RevGAN 22.61 0.816 0.080 21.83 0.769 0.085 25.40 0.862 0.056 23.64 0.844 0.072
MaskGAN 21.40 0.809 0.092 20.41 0.735 0.101 25.24 0.856 0.057 23.15 0.834 0.077
ResViT 22.89 0.822 0.080 21.87 0.786 0.085 25.67 0.872 0.055 24.90 0.873 0.064
SynDiff 21.68 0.814 0.091 20.41 0.755 0.100 25.39 0.872 0.056 24.32 0.867 0.069
ResShift 23.47 0.843 0.075 22.20 0.792 0.082 25.61 0.877 0.055 25.15 0.876 0.062
DFBK (ours) 23.61 0.846 0.074 22.38 0.797 0.080 25.76 0.879 0.054 25.42 0.879 0.060
4.1 Experimental Datasets
For BraTs dataset, we select 40,032 images from 1,251 subjects in the training set with 32 images per subject for model training. Additionally, 7,008 images were chosen from 219 subjects in the validation set with 32 images per subject for generating the target modality and evalution. For ISLES dataset, we select 1,792 images from 28 subjects with 64 images per subject for model training and 2,304 images from 36 subjects with 64 images per subject for evalution. For SynthRAD dataset, we select 9,216 images from 144 subjects with 64 images per subject for model training and 2,304 images from 36 subjects with 64 images per subject for evaluation. The more information on above datasets and training details is available in Supplementary Material A.
4.2 Comparison Approaches
We select three classes of approaches for comparison. The first is GANs-based approaches, EaGAN, RevGAN and MaskGAN. The second is transformer-based approach, ResViT, which is a large model fine-tuning approach. The third is diffusion-based approaches, SynDiff and ResShift. The training parameters are set according to the parameters specified in the original paper, respectively. We use three evaluation metrics, peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and root mean squared error (RMSE), to quantitatively evaluate the generated results to verify the effectiveness of our method.
5 Experimental Results
5.1 Qualitative Results
@@
Fig. 3 shows the results of various methods for medical image translation on BraTs. Ea-GAN introduces grating-like artifacts, and the anatomical structure in some areas is very blurred or even completely distorted as shown in Fig. 3 (a). RevGAN has unrecognizable anatomical structures and introduces additional lesion areas (Fig. 3 (b)). MaskGAN suffers from generating incorrect content and the synthesized images appearing overly blurred (Fig. 3 (c)). ResViT generates a relatively correct anatomical structure, but the brightness of some areas does not match the target modality (Fig. 3 (d)). The anatomical structure generated by SynDiff is incorrect, and the brightness do not match the target modality (Fig. 3 (e)). ResShift generates a generally correct anatomical structure, but the inconsistency of the brightness of certain areas with the target modality is still a noticeable problem (Fig. 3 (f)). In contrast, the images generated by our proposed method demonstrate significantly improved anatomical accuracy and more precise contrast representation. As shown in Figure 3 (g), our approach effectively preserves the structural integrity of anatomical features while maintaining the correct brightness distribution according to target modality. Furthermore, our method achieves the better performance across the quantitative evaluation metrics, underscoring its superiority in medical image translation.
Table 2: Quantitative comparison of different approaches on SynthRAD. The values in the table represent the average values for the entire test dataset. Bold represents the best result and underlined represents the suboptimal result.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
Ea-GAN 24.75 0.852 0.059 20.39 0.679 0.098
RevGAN 26.24 0.872 0.051 22.94 0.798 0.073
MaskGAN 25.56 0.871 0.055 22.04 0.761 0.081
ResViT 26.92 0.887 0.047 22.88 0.801 0.075
SynDiff 26.05 0.877 0.052 20.73 0.705 0.094
ResShift 27.31 0.896 0.045 23.35 0.824 0.071
DFBK (ours) 27.51 0.897 0.044 23.54 0.828 0.069
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
None 25.61 0.877 0.055 25.15 0.876 0.062
+DFB 25.70 0.878 0.055 25.21 0.877 0.062
+KG 25.68 0.877 0.055 25.31 0.878 0.061
+DFB, KG 25.76 0.879 0.055 25.42 0.879 0.060
Fig. 4 shows the results of various methods for medical image translation on ISLES. Ea-GAN cannot generate the correct anatomical structure, and the content is completely distorted, as shown in Figure 4 (a). RevGAN introduces faint raster-like artifacts, and some areas are distorted (Figure 4 (b)). MaskGAN generates blurry content (Figure 4 (c)). The anatomical structure generated by ResViT (Figure 4 (d)) and SynDiff (Figure 4 (e)) exhibits noticeable incorrect, with significant distortions in shape and structure. Additionally, the brightness in certain regions is excessively dark, which fails to align with the target modality. Similarly, the anatomical structure generated by ResShift (Figure 4 (f)) suffers from partial distortions, where some regions exhibit structural inconsistencies. Moreover, the brightness distribution is uneven, with certain areas deviating from the expected intensity levels of the target modality. In contrast, as shown in Figure 4 (g), our method presents more accurate anatomical structures and achieves better metric values.
Fig. 5 shows the results of various methods for medical image translation on SynthRAD. The task of CT to MR translation in SynthRAD is very difficult. Since CT images do not provide information about the internal anatomical structure of the brain, most approaches cannot generate correct or relevant target modality anatomical structures. As shown in Fig. 5 (a), Ea-GAN produces MR images with raster artifacts and lacks clear internal anatomical structures, while its CT images show blurred cranial boundaries. RevGAN can generate relevant MR structures but fails to delineate the lesion area, and its CT images also have blurred cranial boundaries (Fig. 5 (b)). MaskGAN confuses the lesion with normal brain structures in the MR images, and its CT images exhibit incorrect cranial boundaries (Fig. 5 (c)). Although ResViT (Fig. 5 (d)) and SynDiff (Fig. 5 (e)) localize the lesion area in generated MR images, they do not accurately reproduce the overall internal anatomy and CT images generated by ResViT miss part of the cranial boundaries. ResShift produces MR images with generally correct anatomy but still fails to locate the lesion area, and its CT images are also incomplete (Fig. 5 (f)). In contrast, our method accurately localizes the MR lesion area, generates correct anatomical structures, and produces CT images with correct cranial boundaries (Fig. 5 (g)). The more visualization results are in Supplementary Material B.
@@ 5.2 Quantitative Results
Tab. 1 shows the mean values for each quantitative metric, where higher PSNR and SSIM values perform better, while lower RMSE value performs better. Our method outperforms all other approaches in six translation tasks on three datasets. Its superior performance stems from effectively balancing the model’s extraction of high- and low-frequency information and effective prior knowledge guidance, thereby preventing anatomical distortions caused by too fast high-frequency convergence and insufficient low-frequency extraction. In contrast, GAN-based approaches produce poor results due to their inherent instability in training, which makes them particularly susceptible to outliers in low-quality datasets such as BraTs. SynDiff, which incorporates GAN principles into a diffusion model, also shows unsatisfactory performance because its more complex training process is more sensitive to data quality. ResViT achieves relatively good results by leveraging large model fine-tuning and incorporating prior knowledge for cross-domain image synthesis, and ResShift shows suboptimal performance on most metrics through its efficient diffusion process, but the images generated by ResShift still show brightness inconsistency and anatomical distortion due to insufficient low-frequency information extraction. In contrast, our method effectively achieves frequency balance and has better performance.
@@ 5.3 Ablation Study
From Tab. 3, we can see that both dynamic frequency balance and knowledge guidance play an important role in generating the correct anatomical structure and affect most indicators, such as PSNR and SSIM. At the same time, it can be seen that when the source modality and target modality are significantly different, such as the conversion from FLAIR to T1, the knowledge-guided mechanism plays an important role and can effectively guide the model to generate the correct anatomical structure.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
None 25.61 0.877 0.055 25.15 0.876 0.062
+L 25.65 0.878 0.055 25.17 0.876 0,063
+H 25.67 0.877 0.055 25.12 0.874 0,063
+NormAttn 25.59 0.878 0.055 25.09 0.874 0.063
+L, H 25.70 0.878 0.055 25.21 0.877 0.062
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
None 25.61 0.877 0.055 25.15 0.876 0.062
+L 25.66 0.876 0.055 25.21 0.878 0,062
+H 25.59 0.875 0.055 25.15 0.874 0.062
+CrossAttn 25.64 0.877 0.055 25.21 0.875 0.062
+L, H 25.68 0.877 0.055 25.31 0.878 0.061
The components in DFB. In this section, we will verify the impact of low-frequency and high-frequency modules on model performance respectively.
It can be seen from Tab. 4 that only enhancing the low-frequency features (L) can alleviate the poor performance of the model due to too fast high-frequency convergence to a certain extent, while only dynamically balancing the high-frequency features (H) suppresses some high-frequency noise, but the enhancement of the effective high-frequency components actually accelerates the high-frequency convergence of the model and reduces the performance. The improvement of PSNR in the FLAIR to DWI task is related to its similar modality. To further demonstrate the importance of frequency balance, we replace the dynamic frequency balance module with a normal self-attention mechanism (NormAttn). The results show that the performance of the model has slightly decreased. This is because the self-attention mechanism can only enhance global features, but cannot balance high-frequency and low-frequency features, which makes the model further strengthen the convergence of high-frequency features. In contrast, the dynamic frequency balance module can decompose and process frequency components more carefully.
The components in KG. In order to verify the impact of each component of the knowledge-guided mechanism, we conducted an ablation experiment.
As shown in Tab. 5, only fusing the prior knowledge with low-frequency features can guide the model to enhance the extraction of low-frequency features and slightly improve the model performance. However, only fusing the prior knowledge with high-frequency features cannot achieve frequency balance, even if the enhancement of high-frequency features is suppressed. This shows that the enhancement of low-frequency is more important. To further verify the effectiveness of the knowledge-guided mechanism, we replaced it with a cross-attention mechanism (CrossAttn). The results show that the cross-attention mechanism can slightly improve the model performance, but it limits the guidance of the model by the prior knowledge due to the over-enhancement of high-frequency. In contrast, the knowledge-guided mechanism processes high and low frequencies separately, effectively improving model performance.

### v2

# Structure-Accurate Medical Image Translation via Dynamic Frequency Balance and Knowledge Guidance

@@ 1 Introduction
To further improve image generation quality, researchers begin to introduce the prior knowledge of vision language model (VLM) into the generation process [21, 22, 23]. The pre-trained VLM provides rich prior knowledge for image translation, greatly reducing the learning difficulty while ensuring that the generated images have the correct anatomical structure. However, existing algorithms often blindly accept all prior knowledge, resulting in a failure to effectively fuse prior knowledge with the model, which has become a challenge in image translation [24].
@@ 3.2 Medical Image Translation with DFBK
$z^{L}_{s}=S(\frac{Q_{z^{ll}}K_{z^{ll}}^{T}}{\sqrt{d}})V_{z^{ll}},$ (5)
In order to achieve dynamic frequency balance, we use the differential attention mechanism for high-frequency features,
where $\lambda_{q}^{1}$, $\lambda_{q}^{2}$, $\lambda_{k}^{1}$, $\lambda_{k}^{2}$ are learnable parameters, and $\lambda_{init}$ is a constant. Differential attention [40] can dynamically filter out and enhance effective high-frequency information while suppressing useless high-frequency noise.
$z_{s}=IWT(z^{L}_{s},z^{H}_{s}),$ (8)
Figure 3: Comparison of approaches on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) DFBK (ours). The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated image and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa (e.g., blue).
Figure 4: Comparison of approaches on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours). The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively.
$z^{L}_{k}=S(\frac{Q_{z^{ll}}K_{c}^{T}}{\sqrt{d}})V_{c}^{1},$ (9)
Finally, we can obtain the recombination feature $z_{k}$,
$z_{k}=IWT(z^{L}_{k},z^{H}_{k}).$ (11)
Through the dynamic frequency balance module and the knowledge guidance mechanism, we can balance the model’s ability to extract features of high and low frequency information, thereby generating accurate anatomical structures.
@@ 4 Implementation Details
4.1 Experimental Datasets
For BraTs dataset, we select 40,032 images from 1,251 subjects with 32 images per subject for training. Additionally, 7,008 images were chosen from 219 subjects with 32 images per subject for generating the target modality and evalution. For ISLES dataset, we select 1,792 images from 28 subjects with 64 images per subject for training and 2,304 images from 36 subjects with 64 images per subject for evalution. For SynthRAD dataset, we select 9,216 images from 144 subjects with 64 images per subject for training and 2,304 images from 36 subjects with 64 images per subject for evaluation. The more information is available in Supplementary Material A.
4.2 Comparison Approaches
We select three classes of approaches for comparison. The first is GANs-based approaches, EaGAN, RevGAN and MaskGAN. The second is transformer-based approach, ResViT, which is a large model fine-tuning approach. The third is SynDiff whicth is based on diffusion model. The training parameters are set according to the parameters specified in the original paper, respectively. We use two evaluation metrics, peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively evaluate the generated results to verify the effectiveness of our method.
5 Experimental Results
5.1 Qualitative Results
Table 1: Quantitative comparison of different approaches on BraTs and ISLSE. The values in the table represent the average values for the entire test dataset, and higher values are better. Bold indicates the best result, and underlined indicates the suboptimal result.
@@ 5.1 Qualitative Results
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Ea-GAN 22.43 0.825 21.16 0.773 24.18 0.850 21.35 0.810
RevGAN 22.61 0.816 21.83 0.769 25.40 0.862 23.64 0.844
MaskGAN 21.40 0.809 20.41 0.735 25.24 0.856 23.15 0.834
ResViT 22.89 0.822 21.87 0.786 25.67 0.872 24.90 0.873
SynDiff 21.68 0.814 20.41 0.755 25.39 0.872 24.32 0.867
DFBK (ours) 23.61 0.846 22.38 0.797 25.76 0.879 25.42 0.879
@@
Fig. 3 shows the results of various methods for medical image translation on BraTs. Ea-GAN introduces grating-like artifacts, and the anatomical structure in some areas is very blurred or even completely distorted as shown in Fig. 3 (a). RevGAN has unrecognizable anatomical structures and introduces additional lesion areas (Fig. 3 (b)). MaskGAN suffers from generating incorrect content and the synthesized images appearing overly blurred (Fig. 3 (c)). ResViT generates a relatively correct anatomical structure, but the brightness of some areas does not match the target modality (Fig. 3 (d)). The anatomical structure generated by SynDiff is incorrect, and the brightness do not match the target modality (Fig. 3 (e)). In contrast, the images generated by our proposed method demonstrate significantly improved anatomical accuracy and more precise contrast representation. As shown in Figure 3 (f), our approach effectively preserves the structural integrity of anatomical features while maintaining the correct brightness distribution according to target modality. Furthermore, our method achieves the better performance across the quantitative evaluation metrics, underscoring its superiority in medical image translation.
Fig. 4 shows the results of various methods for medical image translation on ISLES. Ea-GAN cannot generate the correct anatomical structure, and the content is completely distorted, as shown in Figure 4 (a). RevGAN introduces faint raster-like artifacts, and some areas are distorted (Figure 4 (b)). MaskGAN generates blurry content (Figure 4 (c)). The anatomical structure generated by ResViT (Figure 4 (d)) and SynDiff (Figure 4 (e)) exhibits noticeable incorrect, with significant distortions in shape and structure. In contrast, as shown in Figure 4 (f), our method presents more accurate anatomical structures and achieves better metric values.
We also shows the results on the difficult task of CT to MR translation on SynthRAD, as shown in Fig. 5 in Supplementary Material B. Since CT images do not provide information about the internal anatomical structure of the brain, most approaches cannot generate correct or relevant target modality anatomical structures. As shown in Fig. 5 (a), Ea-GAN produces MR images with raster artifacts and lacks clear internal anatomical structures, while its CT images show blurred cranial boundaries. RevGAN can generate relevant MR structures but fails to delineate the lesion area, and its CT images also have blurred cranial boundaries (Fig. 5 (b)). MaskGAN confuses the lesion with normal brain structures in the MR images, and its CT images exhibit incorrect cranial boundaries (Fig. 5 (c)). Although ResViT (Fig. 5 (d)) and SynDiff (Fig. 5 (e)) localize the lesion area in generated MR images, they do not accurately reproduce the overall internal anatomy and CT images generated by ResViT miss part of the cranial boundaries. In contrast, our method accurately localizes the MR lesion area, generates correct anatomical structures, and produces CT images with correct cranial boundaries (Fig. 5 (f)). The more visualization results are in Supplementary Material B.
Table 2: Quantitative comparison of different approaches on SynthRAD.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Ea-GAN 24.75 0.852 20.39 0.679
RevGAN 26.24 0.872 22.94 0.798
MaskGAN 25.56 0.871 22.04 0.761
ResViT 26.92 0.887 22.88 0.801
SynDiff 26.05 0.877 20.73 0.705
DFBK (ours) 27.51 0.897 23.54 0.828
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
None 25.61 0.877 25.15 0.876
+DFB 25.70 0.878 25.21 0.877
+KG 25.68 0.877 25.31 0.878
+DFB, KG 25.76 0.879 25.42 0.879
@@ 5.2 Quantitative Results
Tab. 1 and Tab. 3 shows the mean values for each quantitative metric, where higher PSNR and SSIM values perform better. Our method outperforms all other approaches in six translation tasks on three datasets. Its superior performance stems from effectively balancing the model’s extraction of high- and low-frequency information and effective prior knowledge guidance, thereby preventing anatomical distortions caused by too fast high-frequency convergence and insufficient low-frequency extraction. In contrast, GAN-based approaches produce poor results due to their inherent instability in training, which makes them particularly susceptible to outliers in low-quality datasets such as BraTs. SynDiff, which incorporates GAN principles into a diffusion model, also shows unsatisfactory performance because its more complex training process is more sensitive to data quality. ResViT achieves relatively good results by leveraging large model fine-tuning, but still causes anatomical distortion. In contrast, our method effectively achieves frequency balance and has better performance.
@@ 5.3 Ablation Study
From Tab. 3, we can see that both dynamic frequency balance and knowledge guidance play an important role in generating the correct anatomical structure and affect most indicators, such as PSNR and SSIM. At the same time, it can be seen that when the source modality and target modality are significantly different, such as the conversion from FLAIR to T1, the knowledge-guided mechanism plays an important role and can effectively guide the model to generate correct anatomical structures.
The components in DFB. It can be seen from Tab. 5 that only enhancing the low-frequency features (L) can alleviate the poor performance of the model due to too fast high-frequency convergence to a certain extent, while only dynamically balancing the high-frequency features (H) suppresses some high-frequency noise, but the enhancement of the effective high-frequency components actually accelerates the high-frequency convergence of the model and reduces the performance. The improvement of PSNR in the FLAIR to DWI task is related to its similar modality. To further demonstrate the importance of frequency balance, we replace the dynamic frequency balance module with a normal self-attention mechanism (NormAttn). The results show that the performance of the model has slightly decreased. This is because the self-attention mechanism can only enhance global features, but cannot balance high-frequency and low-frequency features, which makes the model further strengthen the convergence of high-frequency features. In contrast, the dynamic frequency balance module can decompose and process frequency components more carefully.
The components in KG. As shown in Tab. 5, only fusing the prior knowledge with low-frequency features can guide the model to enhance the extraction of low-frequency features and slightly improve the model performance. However, only fusing the prior knowledge with high-frequency features cannot achieve frequency balance, even if the enhancement of high-frequency features is suppressed. This shows that the enhancement of low-frequency is more important. To further verify the effectiveness of the knowledge-guided mechanism, we replaced it with a cross-attention mechanism (CrossAttn). The results show that the cross-attention mechanism can slightly improve the model performance, but it limits the guidance of the model by the prior knowledge due to the over-enhancement of high-frequency. In contrast, the knowledge-guided mechanism processes high and low frequencies separately, effectively improving model performance.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
None 25.61 0.877 25.15 0.876
+L 25.65 0.878 25.17 0.876
+H 25.67 0.877 25.12 0.874
+NormAttn 25.59 0.878 25.09 0.874
+L, H 25.70 0.878 25.21 0.877
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
None 25.61 0.877 25.15 0.876
+L 25.66 0.876 25.21 0.878
+H 25.59 0.875 25.15 0.874
+CrossAttn 25.64 0.877 25.21 0.875
+L, H 25.68 0.877 25.31 0.878